Qd = 5 - P , and Qs = -5 + P, then equilibrium price and quantity will be
1
Expert's answer
2019-10-28T11:52:20-0400

"Qd=5-p"

"Qs=-5+p"

To get equilibrium price, equate Qd and Qs and solve for p

"5-p=-5+p"

"10=2p"

"5=p"

Thus, "Ep=5"

To get equilibrium quantity, substitute the equilibrium price to any of the two equations

"Eq=5-p=5-5=0"

"EP=5" and "Eq=0"